OpenStudy (anonymous):

I have a huge problem with my course "Mathematical methods for physicists" Need to solve a problem (in Mathematica for starters), but don't even know, where to begin. Maybe someone could help me out a bit? The problem explained in next post:

OpenStudy (jit4won):

okay

OpenStudy (anonymous):

\[\delta(u)/\delta(t) = a^2*\Delta u + \gamma u + f(M,t)\] \[M=(x,y,z)\]\[x \in R , y \in R , 0<z<H\]\[t >0 , \gamma \ge 0\] Boundry conditions : \[u (t=0) = 0 ; u(z=0)=A ; \delta u / \delta z (z=H) = 0\]\[A \in R\] --- - Give phyisical interpretation (from what i understand, its thermal conduction equation) - find \[v=\lim_{t \rightarrow +\infty} u(M,t) if f=f_0=const\] - What is v if a-> + infinity --- in first line delta before u is Laplacian functions u and v are some temperatures i guess, where v is temperature on boundary.

OpenStudy (anonymous):

correction in first line : \[\delta(u)/\delta(t) = a^2*\Delta u - \gamma u + f(M,t)\]

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

EXACTLY! XD

OpenStudy (anonymous):

im only in algebra one so no way im soving this! Good luck

OpenStudy (anonymous):

Can anyone help me with why the graph isn't plotting??? I'm confused, because everything seems fine to me.

OpenStudy (anonymous):

I would google the subject then mathematica thats what I've done for everything I've done on mathematica they have really awesome guides

OpenStudy (anonymous):

mathematica it*